d a = du I 11. Integrate the following: 3 a) .4 dx b) f(ber - 7 Jahr
Answer #1


b ) 5 1162-32 : Shde = ln (2); sédez urc} 6ſe -- Ja ~ a da da - 6(e) – 7 (In(x)) + c X be In (x) + c f ||(60°-3)dz = (be*- In